# Efficient Learning for Linear Properties of Bounded-Gate Quantum Circuits

### v1

@@ Abstract
The vast and complicated large-qubit state space forbids us to comprehensively capture the dynamics of modern quantum computers via classical simulations or quantum tomography. However, recent progress in quantum learning theory invokes a crucial question: given a quantum circuit containing $d$ tunable RZ gates and $G-d$ Clifford gates, can a learner perform purely classical inference to efficiently predict its linear properties using new classical inputs, after learning from data obtained by incoherently measuring states generated by the same circuit but with different classical inputs? In this work, we prove that the sample complexity scaling linearly in $d$ is necessary and sufficient to achieve a small prediction error, while the corresponding computational complexity may scale exponentially in $d$. Building upon these derived complexity bounds, we further harness the concept of classical shadow and truncated trigonometric expansion to devise a kernel-based learning model capable of trading off prediction error and computational complexity, transitioning from exponential to polynomial scaling in many practical settings. Our results advance two crucial realms in quantum computation: the exploration of quantum algorithms with practical utilities and learning-based quantum system certification. We conduct numerical simulations to validate our proposals across diverse scenarios, encompassing quantum information processing protocols, Hamiltonian simulation, and variational quantum algorithms up to 60 qubits.
I Introduction
@@
Figure 1: Learning protocols for quantum circuits with the bounded number of non-Clifford gates. (a) The visualization of learning bounded-gate quantum circuits with incoherent measurements. Given a circuit composed of $G-d$ Clifford gates and $d$ $\operatorname{RZ}$ gates, a classical learner feeds $n$ classical inputs, i.e., $n$ tuples of the varied angles of $\operatorname{RZ}$ gates, to the quantum device and collects the relevant measured results as data labels. The collected $n$ labeled data are used to train a prediction model $h$ that can perform purely classical inference to predict the linear properties of the generated state over new input $\bm{x}$, i.e., $\operatorname{Tr}(\rho(\bm{x})O)$ with $O$ being an observable sampled from a prior distribution, can be accurately estimated. Following conventions, the interaction between the learner and the system is restrictive in which the learner can only access the quantum computer via incoherent measurements with finite shots. (b) The proposed kernel-based ML model consists of two steps. First, the learner collects the training dataset, i.e., $n$ labeled data via classical shadow. Then, the learner applies shadow estimation and the trigonometric monomial expansion to the collected training dataset to obtain efficient classical representations, where any new input of the explored quantum circuits can be efficiently predicted offline.
Here we answer the above question affirmably by exploring a specific class of quantum circuits that has broad applications in quantum computing. Particularly, this class of $N$-qubit circuits consists of an arbitrary initial state, a bounded-gate circuit composed of $d$ tunable RZ gates and $G-d$ Clifford gates, where the dynamics refer to linear properties of the resulting state measured by the operator $O$. Our first main result uncovers that (i) with high probability, $\widetilde{\Omega}(\frac{(1-\epsilon)d}{\epsilon T})\leq n\leq\widetilde{\mathcal{O}}(\frac{B^{2}d+B^{2}NG}{\epsilon^{2}})$ training examples are sufficient and necessary for a classical learner to achieve an $\epsilon$-prediction error on average, where $B$ refers to the bounded norm of $O$ and $T$ is the number of incoherent measurements; (ii) there exists a class of $G$-bounded-gate quantum circuits that no polynomial runtime algorithms can predict their outputs within an $\epsilon$-prediction error. These results deepen our understanding of the learnability of bounded-gate quantum circuits with classical learners, whereas Ref. [30] focuses on quantum learners. Furthermore, the achieved findings invoke the necessity of developing a learning algorithm capable of addressing the exponential gap between sample efficiency and computational complexity.
@@
II Theoretical Results
We define the state space associated to an $N$-qubit bounded-gate quantum circuit as
@@ II Theoretical Results
[…]
@@
[…]
=$\mathcal{F}=\left\{f(\bm{x},O)=\operatorname{Tr}(\rho(\bm{x})O)\Big{|}\bm{x}\in[-\pi,\pi]^{d}\right\}.$ (3)
Here we suppose $O$ is constituted by multiple local observables with a bounded norm, i.e., $O=\sum_{i=1}^{q}O_{i}$ and $\sum_{l}\|O_{i}\|_{\infty}\leq B$, and the maximum locality of $\{O_{i}\}$ is $K$. This formalism encompasses diverse tasks in quantum computing, e.g., variational quantum algorithms [28], numerous applications of shadow estimation [33], and quantum system certification [29] (see Supplementary Material (SM) A for the elaboration).
The hardness of learning $\mathcal{F}$ in Eq. (3) is explored by separately assessing the required sample complexity and runtime complexity of training a classical learner $h(\bm{x},O)$ that attains a low average prediction error with $f(\bm{x},O)$. Mathematically, the average performance of the learner is expected to satisfy
$\mathbb{E}_{\bm{x}\sim\mathbb{D}_{\mathcal{X}}}\left|h(\bm{x},O)-f(\bm{x},O)\right|^{2}\leq\epsilon,\leavevmode\nobreak\ \forall O\in\mathrm{supp}(\mathbb{D}_{O}),$ (4)
@@
where the classical input $\bm{x}$ is sampled from the distribution $\mathbb{D}_{\mathcal{X}}$ and $\mathrm{supp}$ denotes the support of a probability distribution. Note that our analysis adheres to the conventions of most ML algorithms to train $h(\bm{x},O)$. Namely, as shown in Fig. 1(a), to accommodate the constraints of modern quantum devices, the training data are exclusively gathered by simple measurements [34]. In addition, the classical learner is kept unaware of the circuit layout details, except for the gate count $G$ and the dimension of classical inputs $d$, and can solely access estimated information about quantum states through incoherent measurements with finite shots. Last, the learner $h(\bm{x},O)$ is offline, where the predictions for new inputs $\bm{x}$ is performed solely on the classical side. Given the scarcity of quantum computers, the offline property of $h(\bm{x},O)$ significantly reduces the inference cost for studying incoherent dynamics compared to quantum learners, such as variational quantum algorithms [35, 36].
II.1 Learnability of bounded-gate quantum circuits
@@
The following theorem demonstrates the learnability of $\mathcal{F}$ in Eq. (3), where the formal statement and the proof are presented in SM B-E.
Theorem II.1 (informal).
@@
Following notations in Eq. (3), let $\mathcal{T}=\{\bm{x}^{(i)},\tilde{f}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ be a dataset containing $n$ training examples with $\bm{x}^{(i)}\in[\pi,\pi]^{d}$ and $\tilde{f}_{T}(\bm{x}^{(i)})$ being the estimation of $f(\bm{x}^{(i)},O)$ using $T$ incoherent measurements with $T\leq(N\log 2)/\epsilon$. Then, the training data size
@@
is sufficient and necessary to achieve Eq. (4) with high probability. However, under the assumption that $\mathsf{BQP}\not\subset\mathsf{HeurP}^{\mathbb{D}^{\prime}}/\mathsf{poly}$ for some distribution $\mathbb{D}^{\prime}$, there exists a class of $G$-bounded-gate quantum circuits that no algorithm can achieve Eq. (4) in a polynomial time.
The exponential separation in terms of the sample and computational complexities underscores the non-trivial nature of learning the incoherent dynamics of bounded-gate quantum circuits. While the matched upper and lower bounds in Eq. (5) indicate that a linear number of training examples with $d$ is sufficient and necessary to guarantee a satisfactory prediction accuracy, the derived exponential runtime cost hints that identifying these training examples may be computationally hard. These results enrich quantum learning theory [37, 38], especially for the learnability of bounded-gate quantum circuits. Ref. [30] indicates that for a quantum learner, reconstructing quantum states and unitaries with bounded-gate complexity is sample-efficient but computationally demanding. While Theorem II.1 shows a similar trend, the different learning settings (quantum learner vs. classical learner) and different tasks suggest that the proof techniques in these two studies are inherently different (see SM A for more details). Note that the upper bound does not depend non-trivially on $T$, so we omit it. Besides, an interesting future research direction is to explore novel techniques to match the factors $N$, $B$, and $G$ in the lower and upper bounds, whereas such deviations do not affect our key results.
II.2 A provably efficient protocol to learn bounded-gate quantum circuits
The exponential separation of the sample and computational complexity pinpoints the importance of crafting provably efficient algorithms to learn $\mathcal{F}$ in Eq. (3). To address this issue, here we devise a kernel-based ML protocol adept at balancing the average prediction error $\epsilon$ and the computational complexity, making a transition from exponential to polynomial scaling with $d$ when $\mathbb{D}_{\mathcal{X}}$ is restricted to be the uniform distribution, i.e., $\bm{x}\sim[-\pi,\pi]^{d}$.
@@ II.2 A provably efficient protocol to learn bounded-gate quantum circuits
$\Phi_{\bm{\omega}}(\bm{x})=\prod_{i=1}^{d}\begin{cases}1\leavevmode\nobreak\ &\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=0\\ \cos(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=1\\ \sin(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=-1\end{cases}.$ (8)
A distinctive aspect of our proposal is its capability to predict the incoherent dynamics $\operatorname{Tr}(\rho(\bm{x})O)$ across various $O$ purely on the classical side. This is achieved by storing the shadow information $\{\tilde{\rho}_{T}(\bm{x}^{(i)})\}$ into the classical memory, where the shadow estimation $\{g(\bm{x}^{(i)},O)\}$ for different $\{O\}$ can be efficiently manipulated through classical post-processing (refer to SM G for details).
With the full expansion as $\Lambda=d$, the cardinality of the frequency set $\{\bm{\omega}\}$ in Eq. (7) is $3^{d}$, impeding the computational efficiency of our proposal when the number of $\operatorname{RZ}$ gates becomes large. To remedy this, here we adopt a judicious frequency truncation to strike a balance between prediction error and computational complexity. Define the truncated frequency set as $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\leavevmode\nobreak\ s.t.\leavevmode\nobreak\ \|\bm{\omega}\|_{0}\leq\Lambda\}$. The subsequent theorem provides a provable guarantee for this tradeoff relation, with the formal description and proof deferred to SM F.
Theorem II.2 (Informal).
@@
The obtained results reveal that both sample and computational complexities of $h_{\mathsf{s}}$ are predominantly influenced by the cardinality of $\mathfrak{C}(\Lambda)$, or equivalently the quantity $C$ as $\Lambda=4C/\epsilon$. That is, a polynomial scaling of $|\mathfrak{C}(\Lambda)|$ with $N$ and $d$ can ensure both a polynomial runtime cost to obtain $\kappa_{\Lambda}(\bm{x},\bm{x}^{(i)})g(\bm{x}^{(i)},O)$ and a polynomial sample complexity $n$, leading to an overall polynomial computational complexity of our proposal (see SM G for details). In contrast, for an unbounded $C$ such that $|\mathfrak{C}(\Lambda)|$ exponentially scales with the number of $\operatorname{RZ}$ gates $d$, the computational overhead of $h_{\mathsf{s}}$ becomes prohibitive for a large $d$, aligning with the findings from Theorem II.1.
Figure 3: Numerical results of predicting properties of quantum states evolved by 60-qubit global Hamiltonians. (a) Prediction on the evolved state with $d=1$. The notation $n=a$ refers to the number of training examples used to form the classical representation as $a$. The y-axis denotes the magnetization $\braket{\bar{Z}}=\frac{1}{60}\sum_{i}\braket{Z_{i}}$. (b) Prediction error on the evolved state with $d=30$ with different sizes of the training dataset. The meaning of notations in the box plot is as follows. The label ‘$\Lambda=a$’ stands for that the truncation value is set as $a$. The $x$-axis refers to the varied number of training examples $n$, ranging from $n=10$ to $n=10^{4}$. The $y$-axis refers to the achieved RMS prediction error. All results are obtained using the same random seeds.
We next underscore that in many practical scenarios, the quantity $C$ can be effectively bounded, allowing the proposed ML model to serve as a valuable complement to quantum tomography and classical simulations in comprehending quantum devices. One such scenario involves characterizing near-Clifford quantum circuits consisting of many CI gates and few non-Clifford gates, which find applications in quantum error correction [39, 40] and efficient implementation of approximate t-designs [41]. In this context, adopting the full expansion with $\Lambda=d$ is also computationally efficient, as $|\mathfrak{C}(d)|\sim O(N)$. Meanwhile, when focused on a specific class of quantum circuits described as $\operatorname{\text{CI}}+\operatorname{RZ}$ with a fixed layout, our model surpasses classical simulation methods [42, 43] in runtime complexity by eliminating the dependence on the number of Clifford gates $G-d$.
@@
Last, the complexity bound in Theorem II.2 hints that the locality of the observable $K$ is another factor dominating the performance of $h_{\mathsf{s}}$. This exponential dependence arises from the use of Pauli-based classical shadow, and two potential solutions can be employed to alleviate this influence. The first solution involves adopting advanced variants of classical shadow to enhance the sample complexity bounds [61, 62, 63, 64]. The second solution entails utilizing classical simulators to directly compute the quantity $\{\operatorname{Tr}(\rho(\bm{x}^{(i)})O)\}$ instead of shadow estimation $\{g(\bm{x}^{(i)},O)\}$ in Eq. (6), with the sample complexity summarized in the following corollary.
Corollary II.1 (Informal).
Following notations in Theorem II.2, when $\{\operatorname{Tr}(\rho(\bm{x}^{(i)})O)\}_{i}$ are computed by classical simulators and $n\sim\widetilde{\mathcal{O}}(3^{d}B^{2}d/\epsilon)$, with high probability, the average prediction error is upper bounded by $\epsilon$.
Although using the classical simulators can improve the dependency of the locality of observable and remove the necessity of quantum resources, the price to pay is increasing the computational overhead and only restricting to a small constant $d$. Refer to SM H.1 for the proofs, implementation details, and more discussions.
@@
III Numerical results
We conduct numerical simulations on $60$-qubit quantum circuits to assess the efficacy of the proposed ML model. The omitted details, as well as the demonstration of classically optimizing VQAs with smaller qubit sizes, are provided in SM J.
We first use the proposed ML model to predict the properties of rotational GHZ states. Mathematically, we define an $N$-qubit rotational GHZ states with $N=60$ as |$\ket{\text{GHZ}(\bm{x})}=U(\bm{x})({\ket{0}^{\otimes N}+\ket{1}^{\otimes N}})/{\sqrt{2}}$, where $U(\bm{x})=\operatorname{RY}_{1}(\bm{x}_{1})\otimes\operatorname{RY}_{30}(\bm{x}_{2})\otimes\operatorname{RY}_{60}(\bm{x}_{3})$ and the subscript $j$ refers to apply $\operatorname{RY}$ gate to the $j$-th qubit. At the training stage, we constitute the dataset $\mathcal{T}_{s}$ containing $n=30$ examples, where each example is obtained by uniformly sampling $\bm{x}$ from $[-\pi,\pi]^{3}$ and applying classical shadow to |$\ket{\text{GHZ}(\bm{x})}$ with the shot number $T=1000$.

### v2

@@ ABSTRACT
The vast and complicated large-qubit state space forbids us to comprehensively capture the dynamics of modern quantum computers via classical simulations or quantum tomography. Recent progress in quantum learning theory prompts a crucial question: can linear properties of a large-qubit circuit with $d$ tunable RZ gates and GdG-d Clifford gates be efficiently learned from measurement data generated by varying classical inputs? In this work, we prove that the sample complexity scaling linearly in $d$ is required to achieve a small prediction error, while the corresponding computational complexity may scale exponentially in $d$. To address this challenge, we propose a kernel-based method leveraging classical shadows and truncated trigonometric expansions, enabling a controllable trade-off between prediction accuracy and computational overhead. Our results advance two crucial realms in quantum computation: the exploration of quantum algorithms with practical utilities and learning-based quantum system certification. We conduct numerical simulations to validate our proposals across diverse scenarios, encompassing quantum information processing protocols, Hamiltonian simulation, and variational quantum algorithms up to 60 qubits.
INTRODUCTION
@@
Figure 1: Learning protocols for quantum circuits with the bounded number of non-Clifford gates. (a) The visualization of learning bounded-gate quantum circuits with incoherent measurements. Given a circuit composed of GdG-d Clifford gates and $d$ $\operatorname{RZ}$ gates, a classical learner feeds $n$ classical inputs, i.e., $n$ tuples of the varied angles of $\operatorname{RZ}$ gates, to the quantum device and collects the relevant measured results as data labels. The collected $n$ labeled data are used to train a prediction model $h$ that can perform purely classical inference to predict the linear properties of the generated state over new input $\bm{x}$, i.e., $\operatorname{Tr}(\rho(\bm{x})O)$ with $O$ being an observable sampled from a prior distribution, can be accurately estimated. Following conventions, the interaction between the learner and the system is restrictive in which the learner can only access the quantum computer via incoherent measurements with finite shots. (b) The proposed kernel-based ML model consists of two steps. First, the learner collects the training dataset, i.e., $n$ labeled data via classical shadow. Then, the learner applies shadow estimation and the trigonometric monomial expansion to the collected training dataset to obtain efficient classical representations, where any new input of the explored quantum circuits can be efficiently predicted offline.
Here we answer the above question affirmably by exploring a specific class of quantum circuits that has broad applications in quantum computing. Particularly, this class of $N$-qubit circuits consists of an arbitrary initial state, a bounded-gate circuit composed of $d$ tunable RZ gates and GdG-d Clifford gates, where the dynamics refer to linear properties of the resulting state measured by the operator $O$. Our first main result uncovers that (i) with high probability, $\widetilde{\Omega}(\frac{(1-\epsilon)d}{\epsilon T})\leq n\leq\widetilde{\mathcal{O}}(\frac{B^{2}d+B^{2}NG}{\epsilon^{2}})$ training examples are sufficient and necessary for a classical learner to achieve an $\epsilon$-prediction error on average, where $B$ refers to the bounded norm of $O$ and $T$ is the number of incoherent measurements; (ii) there exists a class of $G$-bounded-gate quantum circuits that no polynomial runtime algorithms can predict their outputs within an $\epsilon$-prediction error. These results deepen our understanding of the learnability of bounded-gate quantum circuits with classical learners, whereas Ref. [30] focuses on quantum learners. Furthermore, the achieved findings invoke the necessity of developing a learning algorithm capable of addressing the exponential gap between sample efficiency and computational complexity.
@@
RESULTS
Problem setup.
— We define the state space associated to an $N$-qubit bounded-gate quantum circuit as
@@ RESULTS
where $\rho_{0}$ denotes an $N$-qubit state and $U(\bm{x})$ refers to the bounded-gate quantum circuit depending on the classical input $\bm{x}$ with $d$ dimensions. Due to the universality of Clifford (abbreviated as $\operatorname{\text{CI}}$) gates with $\operatorname{RZ}$ gates [32], an $N$-qubit $U(\bm{x})$ can always be expressed as
@@
where $d$ is the number of $\operatorname{RZ}$ gates, and $u_{e}$ refers to the unitary composed of $\operatorname{\text{CI}}$ gates and the identity gate $\mathbb{I}_{2}$ with $\operatorname{\text{CI}}=\{H,S,\operatorname{CNOT}\}$. We denote the total number of gates and CI gates in $U(\bm{x})$ as $G$ and GdG-d, respectively. When $\rho(\bm{x})$ is measured by a known observable $O$, its incoherent dynamics are described the mean-value space, i.e., $f(\bm{x},O)=\operatorname{Tr}(\rho(\bm{x})O)$ with $\bm{x}\in[-\pi,\pi]^{d}$. This formalism encompasses diverse tasks in quantum computing, e.g., variational quantum algorithms [28], numerous applications of shadow estimation [33], and quantum system certification [29] (see Supplementary Material (SM) A for the elaboration).
@@
The purpose of most classical learners is to obtain a learning model $h(\bm{x},O)$ to predict $f(\bm{x},O)$. The general paradigm is shown in Fig. 1(a). To accommodate the constraints of modern quantum devices, the training data are exclusively gathered by incoherent measurements. As such, the classical learner feeds $\bm{x}$ into the quantum circuit and receives $\tilde{f}_{T}(\bm{x},O)$, as the estimation of $f(\bm{x},O)$ incurred by $T$ finite shots. By repeating this procedure with varied inputs, the classical learner constructs a training dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{f}_{T}(\bm{x},O)\}_{i=1}^{n}$. Besides, the classical learner is kept unaware of the circuit layout details, except for the gate count $G$ and the dimension of classical inputs $d$. Last, the inferred model $h(\bm{x},O)$ should be offline, where the predictions for new inputs $\bm{x}$ is performed solely on the classical side. Given the scarcity of quantum computers, the offline property of $h(\bm{x},O)$ significantly reduces the inference cost for studying incoherent dynamics compared to quantum learners, such as variational quantum algorithms [34, 35].
According to the above paradigm, the concept class of the mean-value space of bounded-gate circuit is
$\mathcal{F}=\left\{f(\bm{x},O)=\operatorname{Tr}(\rho(\bm{x})O)\Big|U\in\text{Arc}(U,d,G)\right\},$ (3)
where $\bm{x}\in[-\pi,\pi]^{d}$ and $\text{Arc}(U,d,G)$ refers to the set of all possible gate arrangements consisting of $d$ $\operatorname{RZ}$ gates and GdG-d $\operatorname{\text{CI}}$ gates. The hardness of learning $\mathcal{F}$ by classical learners is explored by separately assessing the required sample complexity and runtime complexity of training a classical model $h(\bm{x},O)$ that attains a low average prediction error with $f(\bm{x},O)$. Mathematically, the average performance of the learner is expected to satisfy
$\mathbb{E}_{\bm{x}\sim\mathbb{D}_{\mathcal{X}}}\left|h(\bm{x},O)-f(\bm{x},O)\right|^{2}\leq\epsilon,$ (4)
where the classical input $\bm{x}$ is sampled from the distribution $\mathbb{D}_{\mathcal{X}}$. Here we suppose $O$ consists of $q$ local observables with a $B$-bounded norm, i.e., $O=\sum_{i=1}^{q}O_{i}$ and $\sum_{l}\|O_{i}\|_{\infty}\leq B$, and the maximum locality of $\{O_{i}\}$ is $K$.
@@
Learnability of bounded-gate quantum circuits.— The following theorem demonstrates the learnability of $\mathcal{F}$ in Eq. (3), where the formal statement and the proof are presented in SM B-E (see Methods for the proof sketch).
Theorem 1 (informal).
@@
Following notations in Eq. (3), let $\mathcal{T}=\{\bm{x}^{(i)},\tilde{f}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ be a dataset containing $n$ training examples with $\bm{x}^{(i)}\in[-\pi,\pi]^{d}$ and $\tilde{f}_{T}(\bm{x}^{(i)})$ being the estimation of $f(\bm{x}^{(i)},O)$ using $T$ incoherent measurements with $T\leq(N\log 2)/\epsilon$. Then, the training data size
@@
is sufficient and necessary to achieve Eq. (4) with high probability. However, unless $\mathsf{BQP}\subseteq\mathsf{P/poly}$, there exists a class of $G$-bounded-gate quantum circuits that no algorithm can achieve Eq. (4) in a polynomial time.
The exponential separation in terms of the sample and computational complexities underscores the non-trivial nature of learning the incoherent dynamics of bounded-gate quantum circuits. While the matched upper and lower bounds in Eq. (5) indicate that a linear number of training examples with $d$ is sufficient and necessary to guarantee a satisfactory prediction accuracy, the derived exponential runtime cost hints that identifying these training examples may be computationally hard. Note that the upper bound does not depend non-trivially on $T$, so we omit it. Besides, an interesting future research direction is to explore novel techniques to match the factors $N$, $B$, and $G$ in the lower and upper bounds, whereas such deviations do not affect our key results.
These results enrich quantum learning theory [36, 37], especially for the learnability of bounded-gate quantum circuits. Ref. [30] exhibits that for a quantum learner, reconstructing quantum states and unitaries with bounded-gate complexity is sample-efficient but computationally demanding. While Theorem 1 shows a similar trend, the varied learning settings (quantum learner vs. classical learner) and different tasks suggest that the implications and proof techniques in these two studies are inherently different. As detailed in SM J, these two works are complementary, which collectively reveal the capabilities and limitations of learning bounded-gate quantum circuits.
A provably efficient protocol to learn $\mathcal{F}$.— The exponential separation of the sample and computational complexity pinpoints the importance of crafting provably efficient algorithms to learn $\mathcal{F}$ in Eq. (3). To address this issue, here we devise a kernel-based ML protocol adept at balancing the average prediction error $\epsilon$ and the computational complexity, making a transition from exponential to polynomial scaling with $d$ when $\mathbb{D}_{\mathcal{X}}$ is restricted to be the uniform distribution, i.e., $\bm{x}\sim[-\pi,\pi]^{d}$.
@@
$\Phi_{\bm{\omega}}(\bm{x})=\prod_{i=1}^{d}\begin{cases}1\penalty 10000\ &\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=0\\ \cos(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=1\\ \sin(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=-1\end{cases}.$ (8)
A distinctive aspect of our proposal is its capability to predict the incoherent dynamics $\operatorname{Tr}(\rho(\bm{x})O)$ across various $O$ purely on the classical side. This is achieved by storing the shadow information $\{\tilde{\rho}_{T}(\bm{x}^{(i)})\}$ into the classical memory, where the shadow estimation $\{g(\bm{x}^{(i)},O)\}$ for different $\{O\}$ can be efficiently processed through classical post-processing (refer to Method for details).
With the full expansion as $\Lambda=d$, the cardinality of the frequency set $\{\bm{\omega}\}$ in Eq. (7) is $3^{d}$, impeding the computational efficiency of our proposal when the number of $\operatorname{RZ}$ gates becomes large. To remedy this, here we adopt a judicious frequency truncation to strike a balance between prediction error and computational complexity. Define the truncated frequency set as $\mathfrak{C}(\Lambda)=\{\bm{\omega}|\bm{\omega}\in\{0,\pm 1\}^{d},\penalty 10000\ s.t.\penalty 10000\ \|\bm{\omega}\|_{0}\leq\Lambda\}$. The subsequent theorem provides a provable guarantee for this tradeoff relation, with the formal description and proof deferred to SM F.
Theorem 2 (Informal).
@@
The obtained results reveal that both sample and computational complexities of $h_{\mathsf{s}}$ are predominantly influenced by the cardinality of $\mathfrak{C}(\Lambda)$, or equivalently the quantity $C$ as $\Lambda=4C/\epsilon$. That is, a polynomial scaling of $|\mathfrak{C}(\Lambda)|$ with $N$ and $d$ can ensure both a polynomial runtime cost to obtain $\kappa_{\Lambda}(\bm{x},\bm{x}^{(i)})g(\bm{x}^{(i)},O)$ and a polynomial sample complexity $n$, leading to an overall polynomial computational complexity of our proposal (see Methods and SM G for details). In contrast, for an unbounded $C$ such that $|\mathfrak{C}(\Lambda)|$ exponentially scales with the number of $\operatorname{RZ}$ gates $d$, the computational overhead of $h_{\mathsf{s}}$ becomes prohibitive for a large $d$, aligning with the findings from Theorem 1.
Figure 3: Numerical results of predicting properties of quantum states evolved by 60-qubit global Hamiltonians. (a) Prediction on the evolved state with $d=1$. The notation $n=a$ refers to the number of training examples used to form the classical representation as $a$. The y-axis denotes the magnetization $\braket{\bar{Z}}=\frac{1}{60}\sum_{i}\braket{Z_{i}}$. (b) Pretraining performance of the proposed ML model on a 50-qubit TFIM. The inset presents the prediction error of the proposed ML model for both the ideal (solid bar) and shadowed (Hatching patterns with ‘/’) cases, i.e., $T\rightarrow\infty$ and $T=300$. The main plot depicts the training dynamics of the proposed ML model over $500$ iterations under ideal (circle markers) and shadowed (star markers) settings. The shaded areas represent the variance, while the dashed line indicates the exact ground-state energy.
We next underscore that in many practical scenarios, the quantity $C$ can be effectively bounded, allowing the proposed ML model to serve as a valuable complement to quantum tomography and classical simulations in comprehending quantum devices (See SM K for heuristic evidence). One scenario involves characterizing near-Clifford quantum circuits consisting of many CI gates and few non-Clifford gates, which find applications in quantum error correction [38, 39] and efficient implementation of approximate t-designs [40]. In this context, adopting the full expansion with $\Lambda=d$ is computationally efficient, as $|\mathfrak{C}(d)|\sim O(N)$. Meanwhile, when focused on a specific class of quantum circuits described as $\operatorname{\text{CI}}+\operatorname{RZ}$ with a fixed layout, our model surpasses classical simulation methods [41, 42] in runtime complexity by eliminating the dependence on the number of Clifford gates GdG-d.
@@
Last, the complexity bound in Theorem 2 hints that the locality of the observable $K$ is another factor dominating the performance of $h_{\mathsf{s}}$. This exponential dependence arises from the use of Pauli-based classical shadow, and two potential solutions can be employed to alleviate this influence. The first solution involves adopting advanced variants of classical shadow to enhance the sample complexity bounds [60, 61, 62, 63]. The second solution entails utilizing classical simulators to directly compute the quantity $\{\operatorname{Tr}(\rho(\bm{x}^{(i)})O)\}$ instead of shadow estimation $\{g(\bm{x}^{(i)},O)\}$ in Eq. (157), with the sample complexity summarized in the following corollary.
Corollary 1 (Informal).
Following notations in Theorem 2, when $\{\operatorname{Tr}(\rho(\bm{x}^{(i)})O)\}_{i}$ are computed by classical simulators and $n\sim\widetilde{\mathcal{O}}(3^{d}B^{2}d/\epsilon)$, with high probability, the average prediction error is upper bounded by $\epsilon$.
Although using the classical simulators can improve the dependency of the locality of observable and remove the necessity of quantum resources, the price to pay is increasing the computational overhead and only restricting to a small constant $d$. Refer to SM H for the proofs, implementation details, and more discussions.
Remark. For clarity, we focus on showing how the proposed ML model applied to the bounded-gate circuits with $\operatorname{\text{CI}}$ and $\operatorname{RZ}$ gates. In SM I, we illustrate that our proposal and the relevant theoretical findings can be effectively extended to a broader context, i.e., the circuit composed of $\operatorname{\text{CI}}$ gates alongside parameterized multi-qubit gates generated by arbitrary Pauli strings. In SM J, we show how our proposal differs from Pauli path simulators.
Numerical results.— We conduct numerical simulations on $60$-qubit quantum circuits to assess the efficacy of the proposed ML model. The omitted details, as well as the demonstration of classically optimizing VQAs with smaller qubit sizes, are provided in Method and SM J.